OpenStudy (anonymous):

A projectile is fired straight up at a speed of 13 m/s. How long does it take to reach the top of its motion? The acceleration due to gravity is 9.8 m/s 2 .

OpenStudy (anonymous):

since it is thrown vertically upward, you are only dealing with the vertical components let [up] be positive and [down] as negative initial velocity, vi=13m/s final velocity, Vf=0 m/s --> since you are looking for the time it takes to reach the max height acceleration, a, = -9.8m/s^2 t=? you can use the five kinematics equation to solve this problem \(v_f=v_i +at\) , this equation is more appropriate :)

OpenStudy (anonymous):

evaluating this problem in solving for the variable "t" you'll get \(\large t=\frac{v_f - v_i}{a}\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome :) and \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{e}}\color{#E6E600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#E6E600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{n}}\color{orange}{\text{s}}\color{#E6E600}{\text{t}}\color{green}{\text{u}}\color{blue}{\text{d}}\color{purple}{\text{y}}\color{purple}{\text{!}}\color{red}{\text{ }}\color{orange}{\text{☺}}\color{#E6E600}{\text{}}\end{array} \]